OpenStudy (anonymous):

Help please (: Not good at this, dont understand it Simplify the following expression. 3x − 6y + 2z − 5x − 3y + 8z −2x + 9y + 10z −2x + 9y − 10z −2x − 9y + 10z 2x − 9y + 10z

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

combine all the like terms: 3x - 5x = -6y - 3y = 2z + 8z = string all those answers together with a + and simplify a " + - " to just " - "